A child who weighs 64 lbs has an order for Amoxicillin 200 mg tid. The recommended dose range for this medication is 10-20 mg/kg/day.

  1. What is the child’s recommended daily dose range? __________

  2. What is the individual dose range?______________

  3. Is the order safe?__________________
Answer #1

Weight of the child = 64 lbs = 64 * 0.454 = 29.06 kg

Recommended dose = 29.06 * 10 to 29.06 * 20 = 290.6 mg to 581.2 mg

Individual dose range = 200 mg tid = 200* 3 = 600 mg

No the order is not safe, because it (600 mg)is above the highest recommended dose (581.2 mg) per day.